# CAIMAN: Causal Action Influence Detection for Sample-efficient Loco-manipulation

### v2

@@ 4.1 Low-level Locomotion Policy
The low-level locomotion policy $\bm{\pi}_{l}$ is trained to generate target joint positions $\bm{q}_{j,des}\in\mathbb{R}^{12}$ that track a desired base velocity command $\bm{v}_{r,des}\in\mathbb{R}^{3}$. This command, $\bm{v}_{r,des}=(v^{x}_{r,des},v^{y}_{r,des},\omega^{z}_{r,des})$, specifies linear velocities in the longitudinal ($x$) and lateral ($y$) directions and a yaw rate, all in the robot frame. The desired joint positions $\bm{q}_{j,des}$ are tracked using joint-level proportional-derivative (PD) controllers to produce the corresponding joint torques $\bm{\tau}_{j,des}\in\mathbb{R}^{12}$. The policy is trained with velocity commands sampled uniformly from a predefined range. To ensure robust sim-to-real transfer and hardware performance during contact-rich tasks like pushing, we apply domain randomization and external perturbations following [65]. Further details of the low-level observation $\bm{o}_{l}$ can be found in Appendix A.
4.2 High-level Planning Policy
The high-level policy $\bm{\pi}_{h}$ is trained to generate desired robot velocity commands $\bm{a}_{h}:=\bm{v}_{r,des}$ that achieve successful task completion in object pushing. Its observation $\bm{o}_{h}$ includes the robot’s linear and angular velocities, the poses of the robot and object, the target object position $\bm{p}_{t}=(x_{p},y_{p})\in\mathbb{R}^{2}$, and the previous action—all expressed in the world frame. In scenarios with obstacles (e.g., walls), $\bm{o}_{h}$ also includes their poses in the world frame. For a detailed description of the high-level observation, refer to Appendix A.
We use Proximal Policy Optimization (PPO) [66] as the base RL algorithm, with a simple reward function composed of three terms:
where $\bm{p}_{o}$, $\bm{p}_{t}$ are the current and target object positions, $\epsilon$ is a success threshold, and $\bm{a}_{h}$, $\bm{a}_{h,prev}$ are the current and previous high-level actions, respectively. The exploration bonus $r_{\text{CAI}}$ is derived from the CAI measure $\tilde{C}_{j}$ over the object $j$. We define the CAI reward as:
$r_{\text{CAI}}=\tilde{C}_{j=\text{object}}(s)=\frac{1}{K}\sum_{i=1}^{K}D_{\mathrm{KL}}\left(P_{S_{j}^{\prime}\mid S=s,A=a^{(i)}}\bigg{\|}\frac{1}{K}\sum_{k=1}^{K}P_{S_{j}^{\prime}\mid S=s,A=a^{(k)}}\right),$ (2)
given $K$ actions $\{a^{(i)}\}_{i=1}^{K}$ sampled from the policy. This approximation estimates the marginal $P_{S_{j}^{\prime}\mid s}$ using Monte Carlo sampling. We model the transition distribution $P_{S_{j}^{\prime}\mid S=s,A=a}$ as a fixed-variance Gaussian (details in Section 4.3), which enables closed-form KL divergence calculation using the Gaussian mixture approximation [67].
Finally, we foster more directed exploration by injecting time-correlated noise into the action sampling process during training [68, 69]. For more details on training the high-level policy, we refer the reader to Appendix D.
4.3 Dynamics learning
$\tilde{\bm{p}}^{\prime}_{o}=\begin{cases}\bm{p}_{o}+\delta t\cdot(\bm{a}_{h,xy}\cdot\delta\hat{\bm{p}})\cdot\delta\hat{\bm{p}},\quad&\text{if }\left\lVert\bm{p}_{o}-\bm{p}_{r}\right\rVert_{2}\leq\epsilon_{p}\text{ and }\bm{a}_{h,xy}\cdot\delta\hat{\bm{p}}>0\\ \bm{p}_{o},&\text{otherwise.}\end{cases}$ (4)
Although the high-level action $\bm{a}_{h}$ could, in principle, be sampled from the full support of the desired velocity $\bm{v}_{r,des}$ [26], not all commands are feasible for the low-level controller to execute under the robot’s current state. Therefore, we define $\bm{a}_{h}=\tilde{\bm{v}}_{r}$ to be the achievable velocity, and use $\tilde{\bm{v}}_{r}$ for both dynamics learning and CAI computation. Given training samples $\mathcal{D}={(\bm{s}_{h}^{(i)},\bm{a}_{h}^{(i)},\bm{s}^{\prime(i)}_{h})}$, we have access to the achieved robot velocity $\bm{v}^{\prime}_{r}$ at the next timestep. When computing CAI as an exploration reward, we assume that the robot’s velocity can only change within a limited range over one high-level step. Thus, for any state $\bm{s}_{h}$, we sample the action $\bm{a}_{h}$ from a bounded range centered on the current velocity:
$\bm{a}_{h}=\tilde{\bm{v}}_{r}\sim\mathcal{U}[\bm{v}_{r}\pm\delta\bm{v}_{r}]$ (7)
where $\delta\bm{v}_{r}$ is a fixed velocity deviation range. Detailed values for the velocity limits $(\delta v_{r}^{x},\delta v_{r}^{y},\delta\omega_{r}^{z})$ are provided in Appendix D.
5 Experiments
Figure 3: Illustrations for the Single-object (left), Single-wall (middle), and Multi-wall (right) tasks. The yellow sphere denotes the object’s target position.
We train both the high-level and low-level policies using Isaac Sim [70]. Training hyperparameters and additional details are provided in Appendix D. We evaluate CAIMAN on three pushing tasks of increasing difficulty, illustrated in Figure 3. For the Single-object task, the robot must manipulate a single cuboid object to a fixed target position. The Single-wall task includes a fixed wall that blocks the direct path between the object and the target. The robot must navigate the object around the wall to reach the goal. Finally, the Multi-wall task includes multiple fixed walls that obstruct the path between the object and the target. The robot must maneuver the object through these obstacles to complete the task. The presence of fixed wall obstacles in the single-wall and multi-wall tasks introduces regions in the state space where object dynamics become more complex, reducing the robot’s ability to exert consistent influence over the object. For all tasks, we train with a fixed target position. A full description of scene configurations is provided in Appendix B.
Baselines
We compare our approach against several baselines. The Heuristics baseline trains the high-level policy using a distance-based heuristic reward, $r_{\text{heu}}=\exp\left(-\left\lVert\bm{p}_{r}-\bm{p}_{o}\right\rVert_{2}\right)$, encouraging the robot to minimize its distance to the object. This mirrors the main explorative reward from prior work [25]. We also conduct comparisons with other intrinsic motivation algorithms, namely Random Network Distillation (RND) [71]. We implement two variants: RND-full, which mirrors the original RND implementation and computes the prediction error for the full state, and RND-object, which computes the prediction error for the object’s state only. As for ablations, we implement CAI-kinematic, where CAI is computed using only the kinematic prior (Equation 4), and and CAI-learned, where CAI is computed using a fully learned dynamics model (no prior).
For detailed formulations of all reward functions and their coefficients, see Appendix C.
5.1 Simulation results
HeuristicsCAI-kinematicCAI-learnedRND-fullRND-objectCAIMAN
We evaluate the learning performance of each method by measuring the success rate for each task. A task is considered successfully completed if the distance between the object and the target is below a threshold $\epsilon_{s}$ at the end of an episode. The results are shown in Figure 4.
With only sparse task rewards, task-relevant skill acquisition heavily relies on effective exploration. In the Single-object task, all methods—except RND-full—achieve some degree of success. CAIMAN, RND-object, and Heuristics perform best, while CAI-kinematic and CAI-learned underperform. The low performance of CAI-learned is likely due to the high data and time demands required to learn an accurate dynamics model from scratch. In contrast, CAI-kinematic performs comparably to the top methods, suggesting that the naive kinematic prior is sufficient for simpler tasks with straightforward object interactions.
In the more complex Single-wall and Multi-wall tasks, only CAIMAN and RND-object demonstrate notable success, with RND-object exhibiting lower asymptotic performance than CAIMAN in both scenarios. This highlights CAIMAN’s unique ability to guide task-relevant exploration even in the presence of obstacles, validating its effectiveness in cluttered environments without relying on dense extrinsic rewards. The failure of CAI-kinematic and CAI-learned in these tasks emphasizes the importance of combining a kinematic prior with a learned residual model to efficiently capture the complex object dynamics needed for effective exploration.
RND-full results in agents randomly exploring the state space without yielding meaningful behavior in all tasks. This finding reinforces the need for object-centric exploration bonuses when the skill to be learned involves physical interaction with objects. The performance gap between RND-object and CAIMAN can be partially attributed to the phenomenon of detachment [72], where the agent drifts away from reward-depleted areas and continuously pursues new regions with higher intrinsic reward, potentially losing avenues of task-relevant exploration.
5.2 Leveraging pretrained dynamics for new tasks
Figure 5: Single wall task with target positions randomly sampled within an area in front of the wall.
Figure 6: Learning the single wall random target task using learning from scratch and learning with models transferred from the fixed target task.
Since our framework learns the environment dynamics, we propose that the dynamics residual obtained from a previous training can be reused across different tasks—as long as the underlying dynamics remain consistent. We hypothesize that leveraging a pretrained model improves the accuracy of CAI estimates in the early stages of training, thereby enhancing the exploration signal and significantly boosting sample efficiency.
To test this hypothesis, we consider a generalization of the Single-wall task in which the target position is randomized instead of fixed, sampled uniformly within a predefined area as illustrated in Figure 6. Detailed scene configurations are provided in Appendix B. We compare the original CAIMAN, which learns the dynamics residual from scratch, to one that reuses a pretrained residual model. For the latter, we initialize the residual with one from the original Single-wall task and continuously fine-tune it during training with data collected from the new generalized task.
5.3 Hardware deployment
Figure 7: Snapshots from the hardware deployment for the single object and single wall tasks. The mass of the box is 5.5 kilograms with a dimension of $(0.55,0.55,0.5)$ meters.
We validated our trained policies on real-world quadruped pushing tasks with Unitree Go2, as shown in Figure 7. To bridge the sim-to-real gap, we applied domain randomization to object mass and friction during high-level policy training. Additional details are provided in Appendix E. All entities in the environment were tracked using an external motion capture system. The trained policies are deployed directly to the robot and successfully execute pushing tasks without requiring additional fine-tuning. We refer interested readers to the supplementary video for more details.
6 Conclusion
We present CAIMAN, a general framework for training whole-body pushing skills in legged robots. CAIMAN adopts a hierarchical control strategy that decouples locomotion from high-level planning. By introducing an intrinsic exploration bonus based on CAI, our method encourages the robot to gain control over relevant entities in its environment under sparse task signals—eliminating the need for tedious, hand-crafted reward shaping, even in cluttered scenarios with obstacles. To further enhance training efficiency, we bootstrap the CAI computation using a simple yet effective kinematic prior, and refine it with a learned residual dynamics model. We evaluate CAIMAN on a suite of quadruped pushing tasks and show that it consistently outperforms competitive baselines in terms of sample efficiency. We also demonstrate that the learned dynamics residual can be reused to accelerate training for new tasks, highlighting the transferability of our approach. Finally, we validate our method on real hardware, seamlessly deploying the trained policies without the need for fine-tuning. Overall, CAIMAN offers a robust and scalable solution for learning physically grounded manipulation behaviors in legged robots, paving the way toward more autonomous and adaptable robotic systems.
7 Limitations
Although CAI-based exploration provides meaningful guidance for robot-object interaction, it inevitably relies heavily on an accurate dynamics model. Currently, CAIMAN trains the high-level policy alongside the dynamics residual, which may lead to undesired exploratory behaviors in the early stages of training when the dynamics model is still inaccurate. This can be mitigated by leveraging a pretrained world model. Our deployment pipeline currently relies on an external motion capture system. We leave the incorporation of an onboard perceptive module for future work. Another potential direction for the future is to extend CAIMAN to enable exploration for multi-object interactions.
Acknowledgments
We thank Dongho Kang, Taerim Yoon, Zijun Hui, Flavio De Vincenti and Hélène Stefanelli for developing the hardware testing framework.
Appendix A Detailed Observation Space
Below we list the observation space for low-level, high-level policies and the high-level state for calculating CAI.
@@ Appendix A Detailed Observation Space
$\bm{v}_{r,des}=(v^{x}_{r,des},v^{y}_{r,des},\omega^{z}_{r,des})\in\mathbb{R}^{3}$ Desired velocity command
@@
Table A.1: Detailed observation space for each module.
Appendix B Scene Configurations
The detailed initial pose for each entity and target position are shown in Table B.1. In simulation, the object measures $0.5$ meters in length, width, and height, and weighs $5.0$ kilograms. The wall in the single-wall task has dimensions $(0.1,1.0,0.5)$ meters, while walls in the multi-wall task measure $(0.1,1.25,0.5)$ meters. For the transfer task of Single-wall, we sample the target position within the range of ($\mathcal{U}[3.0,4.0]$, $\mathcal{U}[-1.0,1.0]$).
Task Single-object Single-wall Multi-wall
Robot (0.0, 0.0, $\mathcal{U}(-\pi,\pi]$) (0.0, 0.0, $\mathcal{U}(-\pi,\pi]$) (0.0, 0.0, $\mathcal{U}(-\pi,\pi]$)
Object (1.0, 0.0, 0.0) (1.0, 0.0, 0.0) (1.0, 0.0, 0.0)
Target (3.0, 0.0) (3.5, 0.0) (4.0, 0.0)
Wall 1 - (2.0, 0.0, 0.0) (2.0, 0.0, 0.0)
Wall 2 - - (3.25, 1.25, 0.0)
Wall 3 - - (3.25, -1.25, 0.0)
Table B.1: We list the scene configurations, where the initial pose $(x,y,\psi)$ for every existing entity and the target position $(x,y)$ for each task (Single-object, Single-wall, Multi-wall) is shown.
Appendix C Rewards and Coefficients
We list the reward functions and corresponding reward parameters for all tasks and training methods in Table C.1. The RND reward term is defined as $r_{RND}=\lVert f(s)-\hat{f}(s)\rVert_{2}$, where $s$ is the state to be encoded, and $f(s)$ and $\hat{f}(s)$ are the target and predictor networks, respectively.
@@ Appendix C Rewards and Coefficients
Table C.1: Simulation training reward parameters. All CAI methods (CAI-kinematic, CAI-learned, CAIMAN) use the CAI reward function, all RND methods (RND-full, RND-object) use the RND reward function, while only the Heuristics method uses the heuristics reward function.
Appendix D Additional training details
The low-level policy operates with a control interval of 0.02 seconds, while the high-level policy has a control interval of 0.2 seconds. For every iteration through the high-level control loop, we step the environment by 0.2 seconds, compute the rewards and terminations, collect transitions for the high-level and dynamics policies, and calculate and add the CAI exploration bonus to the rewards buffer. We trained 1500 iterations for the single object and single wall tasks and 2000 iterations for the multi-wall task, where each iteration consists of 10 high-level control steps. All tasks have an episode length of 20 seconds.
In addition to CAI, we also generate meaningful exploration through colored noise during policy rolling out [68, 69]. By sampling from time-correlated actions, we reduce the possibility of meaningless back-and-forth behavior that could result from commonly used white-noise samples. We select a correlation strength parameterized as $\beta=0.5$, corresponding to a colored noise between white and pink.
We list all the hyperparameters used for training the high-level loco-manipulation policy and dynamics residual model in Table D.1.
Environment hyperparameters
number of environments 4096
$v_{r,des}^{x}$ (m/s) $[-1.0,1.0]$
$v_{r,des}^{y}$ (m/s) $[-1.0,1.0]$
$\omega_{r,des}^{z}$ (rad/s) $[-1.0,1.0]$
$\delta{v}_{r}^{x}$ (m/s) $0.3$
$\delta{v}_{r}^{y}$ (m/s) $0.3$
$\delta{\omega}_{r}^{z}$ (rad/s) $0.4$
threshold for kinematic prior model $\epsilon_{p}$ (m) $0.7$
threshold for sparse task reward $\epsilon$ (m) $0.1$
success rate threshold $\epsilon_{s}$ (m) $0.15$
high-level policy frequency (Hz) $5$
low-level policy frequency (Hz) $50$
number of sampled action in CAI $K$ 64
PPO hyperparameters
policy network $[512,256,128]$
policy activation function ELU
policy initial standard deviation $1.0$
value network $[512,256,128]$
value activation function ELU
number of mini-batches $4$
number of epochs per iteration $5$
clip threshold $0.2$
learning rate schedule adaptive
desired KL divergence $0.01$
value function loss coefficient $1.0$
entropy coefficient $0.005$
max gradient norm $1.0$
$\gamma$ $0.99$
$\lambda$ $0.95$
Dynamics hyperparameters
fixed variance $\sigma$ $1.0$
residual model $[256,128]$
batch size $4096$
number of epochs per iteration $8$
learning rate $0.0001$
buffer size $10000000$
Table D.1: High-level policy and dynamics residual training hyperparameters.
Appendix E Domain randomization for hardware deployment
Table E.1 presents the environment parameters that were randomized in simulation during training policies for hardware deployment.
Parameter Range
Object mass $(kg)$ $\mathcal{U}[3.5,7.5]$
Object friction coefficient $\mathcal{U}[0.5,1.5]$
Initial robot position $(m)$ x: $\mathcal{U}[-0.1,0.1]$, y: $\mathcal{U}[-0.1,0.1]$
Initial object position $(m)$ x: $\mathcal{U}[0.9,1.1]$, y: $\mathcal{U}[-0.1,0.1]$
Table E.1: Randomized parameters in simulation for training HL policies for hardware.
Appendix F Additional Experiment: Dense task reward
Though we primarily focused on a sparse task reward, we also trained all tasks in simulation under a dense task reward for all methods. The dense task reward, defined as $r_{\text{task}}^{dense}=\exp{-\left\lVert\bm{p}_{o}-\bm{p}_{t}\right\rVert_{2}}$, awards an amount inversely proportional to the euclidean distance between the object and the target. Combining the dense task rewards with a distance-based heuristic reward resembles the settings from prior work [25].
Task performance during training is shown in Figure F.1. For learning the single object task with dense rewards, all methods demonstrate similar learning speeds and success rates, which indicates that within simpler, less cluttered environments, a dense reward scheme sufficiently guides task learning.
However, for the dense reward single wall and multi-wall tasks, all CAI-imbued learning methods (CAIMAN, CAI-kinematic, CAI-learned) exhibit better sample efficiency over the heuristics method, indicating that learning the actions necessary for obstacle navigation is facilitated through CAI exploration. CAIMAN and CAI-kinematic achieve the fastest convergence and the highest success rates, while CAI-learned is slightly less sample efficient. This can be attributed to the fact that learning dynamics from scratch initially produces erroneous predictions, which results in inaccurately high CAI scores and diverts the robot from acquiring task specific skills. In contrast, with only the kinematic prior model, the resulting exploration behavior is less adversely impacted when skill learning is heavily guided by dense rewards.
HeuristicsCAI-kinematicCAI-learnedCAIMAN
Figure F.1: Policy success rate during training under a dense task reward for all methods. All experiments utilize a minimum of 2 seeds.
Appendix G Additional Experiment: Cart
To demonstrate CAIMAN’s effectiveness in handling complex objects with asymmetric dynamics, we provide an additional experiment: pushing a simplified cart with two normal and two caster wheels, as shown in Figure G.1. The scene configuration is identical to that of the Single object task, with the box being replaced by the cart. Our results demonstrate that CAIMAN is more sample efficient and achieves a higher asymptotic performance (approx. 90% success rate) than most baselines and is on par with the best competitor. Notably, this experiment did not require any changes to the framework or physics prior, highlighting the capability of the learned model in estimating complex dynamics and the effectiveness of CAIMAN for learning with irregular objects. This environment has no obstacles, but we hypothesize that cluttered settings with walls would further emphasize CAIMAN’s advantages over baselines.

### v3

@@ IV-A Low-level Locomotion Policy
The low-level locomotion policy $\bm{\pi}_{l}$ is trained to generate target joint positions $\bm{q}_{j,des}\in\mathbb{R}^{12}$ that track a desired base velocity command $\bm{v}_{r,des}\in\mathbb{R}^{3}$. This command, $\bm{v}_{r,des}=(v^{x}_{r,des},v^{y}_{r,des},\omega^{z}_{r,des})$, specifies linear velocities in the longitudinal ($x$) and lateral ($y$) directions and a yaw rate, all in the robot frame. The desired joint positions $\bm{q}_{j,des}$ are tracked using joint-level proportional-derivative (PD) controllers to produce the corresponding joint torques $\bm{\tau}_{j,des}\in\mathbb{R}^{12}$. The policy is trained with velocity commands sampled uniformly from a predefined range. To ensure robust sim-to-real transfer and hardware performance during contact-rich tasks like pushing, we apply domain randomization and external perturbations following [rudin2022learning]. The complete low-level observation space $\bm{o}_{l}$ is shown in Table I.
@@
$\bm{v}_{r,des}\in\mathbb{R}^{3}$ Desired velocity command
@@
TABLE I: Detailed observation space for each module.
IV-B High-level Planning Policy
The high-level policy $\bm{\pi}_{h}$ is trained to generate desired robot velocity commands $\bm{a}_{h}:=\bm{v}_{r,des}$ that achieve successful task completion in object pushing. Its observation $\bm{o}_{h}$ includes the robot’s linear and angular velocities, the poses of the robot and object, the target object position $\bm{p}_{t}=(x_{p},y_{p})\in\mathbb{R}^{2}$, and the previous action—all expressed in the world frame. In scenarios with obstacles (e.g., walls), $\bm{o}_{h}$ also includes their poses in the world frame. The complete high-level observation space $\bm{o}_{h}$ is shown in Table I.
We use Proximal Policy Optimization (PPO) [schulman2017proximal] as the base RL algorithm, with a simple reward function composed of three terms:
where $\bm{p}_{o}$, $\bm{p}_{t}$ are the current and target object positions, $\epsilon$ is a success threshold, and $\bm{a}_{h}$, $\bm{a}_{h,prev}$ are the current and previous high-level actions, respectively. The exploration bonus $r_{\text{CAI}}$ is derived from the CAI measure $C^{j}$ from (1), where we choose the object to be the entity of interest $j$. Resorting to an approximation $\tilde{C}_{j}$, we compute the reward $r_{\text{CAI}}$ at a given state $S=\bm{s}$ as
$r_{\text{CAI}}=\tilde{C}_{j=\text{object}}(s)=$ (3)
$\frac{1}{K}\sum_{i=1}^{K}D_{\mathrm{KL}}\left(P_{S_{j}^{\prime}\mid S=s,A=a^{(i)}}\bigg\|\frac{1}{K}\sum_{k=1}^{K}P_{S_{j}^{\prime}\mid S=s,A=a^{(k)}}\right),$
given $K$ actions $\{a^{(i)}\}_{i=1}^{K}$ sampled from the policy. In this work, we find it sufficient to use $K=64$. This approximation estimates the marginal $P_{S_{j}^{\prime}\mid s}$ using Monte Carlo sampling. We model the transition distribution $P_{S_{j}^{\prime}\mid S=s,A=a}$ as a fixed-variance Gaussian (details in Section IV-C), which enables closed-form KL divergence calculation using the Gaussian mixture approximation [durrieu2012lower].
Finally, we foster more directed exploration by injecting time-correlated noise into the action sampling process during training [eberhard2023pink, hollenstein2024colored]. Sampling from time-correlated actions reduces the possibility of meaningless back-and-forth behavior that could result from commonly used white-noise samples. We select a correlation strength parameterized as $\beta=0.5$, corresponding to a colored noise between white and pink.
IV-C Dynamics learning
[…]
Although the high-level action $\bm{a}_{h}$ could, in principle, be sampled from the full support of the desired velocity $\bm{v}_{r,des}$ [seitzer2021causal], not all commands are feasible for the low-level controller to execute under the robot’s current state. Therefore, we define $\bm{a}_{h}=\tilde{\bm{v}}_{r}$ to be the achievable velocity, and use $\tilde{\bm{v}}_{r}$ for both dynamics learning and CAI computation. Given training samples $\mathcal{D}={(\bm{s}_{h}^{(i)},\bm{a}_{h}^{(i)},\bm{s}^{\prime(i)}_{h})}$, we have access to the achieved robot velocity $\bm{v}^{\prime}_{r}$ at the next timestep. When computing CAI as an exploration reward, we assume that the robot’s velocity can only change within a limited range over one high-level step. Thus, for any state $\bm{s}_{h}$, we sample the action $\bm{a}_{h}$ from a bounded range centered on the current velocity:
$\bm{a}_{h}=\tilde{\bm{v}}_{r}\sim\mathcal{U}[\bm{v}_{r}\pm\delta\bm{v}_{r}]$ (8)
where $\delta\bm{v}_{r}$ is a fixed range of velocity deviation. The limits were empirically determined to be $(\delta v_{r}^{x},\delta v_{r}^{y},\delta\omega_{r}^{z})=(0.3,0.3,0.4)$.
Figure 4: The Single Object (left), Single Wall (middle), and Multi-Wall (right) tasks. The yellow sphere denotes the object’s target position.
V EXPERIMENTS
We trained both high- and low-level policies using Isaac Sim [mittal2023orbit]. The low-level policy operates with a control interval of 0.02 seconds, while the high-level policy has a control interval of 0.2 seconds. At each iteration of the high-level control loop, we advance the environment by 0.2 seconds, compute rewards and terminations, collect transitions for both the high-level and dynamics policies, and augment the rewards buffer with the CAI exploration bonus.
We evaluated CAIMAN on three pushing tasks of increasing difficulty, illustrated in Figure 4. In the Single-object task, the robot pushes a single cuboid object to a fixed target position. The Single-wall task introduces a fixed wall that blocks the direct path to the target, requiring the robot to maneuver the object around the obstacle. The Multi-wall task further increases complexity with multiple fixed walls through which the robot must navigate the object. The wall obstacles in the single-wall and multi-wall tasks introduce regions in the state space where the object’s dynamics become more complex, thereby reducing the robot’s ability to exert consistent influence over the object. We trained 1500 iterations for the single object and single wall tasks, and 2000 iterations for the multi-wall task, where each iteration consists of 10 high-level control steps. All tasks have an episode duration of 20 seconds and use a fixed target position.
@@ V EXPERIMENTS
TABLE II: Simulation training reward parameters. All CAI methods (CAI-kinematic, CAI-learned, CAIMAN) use the CAI reward function, all RND methods (RND-full, RND-object) use the RND reward function, while only the Heuristics method uses the heuristics reward function.
We compared our approach against several baselines. The Heuristics baseline trains the high-level policy using a distance-based heuristic reward, $r_{\text{heu}}=\exp\left(-\left\lVert\bm{p}_{r}-\bm{p}_{o}\right\rVert_{2}\right)$, which encourages the robot to minimize its distance to the object and mirrors the primary exploration reward used in prior work [jeon2023learning]. We further evaluated CAIMAN against intrinsic motivation algorithms, specifically Random Network Distillation (RND) [burda2018exploration], implementing two variants: RND-full, which follows the original formulation and computes prediction error over the full state, and RND-object, which restricts prediction error to the object state only. For ablation studies, we introduced CAI-kinematic, where CAI is computed using only the kinematic prior (Equation 5), and CAI-learned, where CAI relies on a fully learned dynamics model (no prior). The reward functions and corresponding parameters for all tasks and training methods are summarized in Table II. The RND reward term is given as $r_{RND}=\lVert f(s)-\hat{f}(s)\rVert_{2}$, where $s$ is the state to be encoded, and $f(s)$ and $\hat{f}(s)$ denote the target and predictor networks, respectively.
@@
V-A Simulation results
We evaluated the learning performance of each method by measuring the success rate for each task. A task is successfully completed if the distance between the object and the target is below a threshold $\epsilon_{s}$ at the end of an episode. The results are shown in Figure 5.
With only sparse task rewards, task-relevant skill acquisition heavily relies on effective exploration. In the Single-object task, all methods—except RND-full—achieve some degree of success. CAIMAN, RND-object, and Heuristics perform the best, while CAI-kinematic and CAI-learned underperform. The low performance of CAI-learned is likely due to the high data and time demands required to learn an accurate dynamics model from scratch. In contrast, CAI-kinematic performs comparably with the top methods, suggesting that the naive kinematic prior is sufficient for simpler tasks with straightforward object interactions. In the more complex Single-wall and Multi-wall tasks, only CAIMAN and RND-object demonstrate notable success, with RND-object showing lower asymptotic performance than CAIMAN in both scenarios. This highlights CAIMAN’s unique ability to guide task-relevant exploration even in the presence of obstacles, validating its effectiveness in cluttered environments and independence from dense extrinsic rewards. The failure of CAI-kinematic and CAI-learned in these tasks emphasizes the importance of combining a kinematic prior with a learned residual model to efficiently capture the complex object dynamics needed for effective exploration.
RND-full results in agents randomly exploring the state space without yielding meaningful behavior in all tasks. This finding reinforces the need for object-centric exploration bonuses when the skill to be learned involves physical interaction with objects. The performance gap between RND-object and CAIMAN can be partially attributed to the phenomenon of detachment [ecoffet2019go], where the agent drifts away from reward-depleted areas and continuously pursues new regions with higher intrinsic reward, potentially losing avenues of task-relevant exploration.
To validate CAI as an effective exploration signal, Figure 6 explicitly plots the raw CAI reward alongside robot-object interactions in the Single-object task. Elevated CAI values during contact states empirically confirm the role of CAI in guiding exploration.
CAI rewardcontact
Figure 6: Scaled CAI reward overlaid with a binary indicator of robot–object contact (1-contact, 0-no contact) over one episode of the Single-object task. Representative environment frames are shown above.
V-B Loco-manipulation of irregular objects
To demonstrate CAIMAN’s effectiveness in handling complex objects with asymmetric dynamics, we conducted an additional experiment using a simplified cart with two normal and two caster wheels, as illustrated in Figure 7. The scene configuration is identical to that of the Single-object task, with the cuboid being replaced by the cart. Our results show that CAIMAN is more sample efficient and achieves a higher asymptotic performance (approx. 90% success rate) than most baselines and is on par with the best competitor. In particular, this experiment did not require any changes to the original framework or to the kinematics prior, which was derived assuming an object with symmetric motion. These results highlight the ability of the learned model to accurately capture complex dynamics and confirm CAIMAN’s capability in learning with irregular objects. This environment has no obstacles, but we hypothesize that adding walls would further emphasize CAIMAN’s advantages over other baselines.
HeuristicsCAI-kinematicCAI-learnedRND-fullRND-objectCAIMAN